OpenStudy (anonymous):

Find the inverse function: y=-2cos(3x)

OpenStudy (anonymous):

The inverse as in it's negative not it's opposite

OpenStudy (rizags):

the inverse is its reflection over y=x

OpenStudy (rizags):

replace x in the initial equation, with y, and thats your inverse!

OpenStudy (anonymous):

I got that far then do I divide the 3 or divide the -2 first? @Rizags

OpenStudy (rizags):

divide by -2

OpenStudy (rizags):

tell me what you get after didividing by -2

OpenStudy (anonymous):

x/-2=cos3y

OpenStudy (rizags):

now, take the INVERSE cos of each side. Do you know what that is?

OpenStudy (anonymous):

Nope

OpenStudy (rizags):

well, the INVERSE cos essentially undoes the cosine. after taking the inverse cos, you get \[\cos^{-1} (\frac{ -x }{ 2 })=3y\]

OpenStudy (rizags):

then divide both sides by three

OpenStudy (rizags):

tell me when you've done that

OpenStudy (anonymous):

x/-2/3?cos^-1=y

OpenStudy (anonymous):

@Rizags

OpenStudy (rizags):

lemme write it out

OpenStudy (rizags):

\[y = ±(\frac{1}{3} \cos^{-1}(\frac {-x}{2})\]

OpenStudy (rizags):

thats it

OpenStudy (anonymous):

If I'm solving for cosx to find the domain, how would I go about doing that? @Rizags

OpenStudy (rizags):

the domain is all reals

OpenStudy (perl):

your inverse function is $$ \Large y = \frac{1}{3} \cos^{-1}(\frac {-x}{2}) $$ the domain of y = arcos (x) is [-1,1] so the domain of arcos(-x/2) is -1 <= -x/2 <= 1

OpenStudy (perl):

that gives you -2 <= x <= 2 for your domain